OpenStudy (anonymous):

Add and simplify if possible:

OpenStudy (anonymous):

\[x+3 \over x-4 \] + \[x \over 4-x\]

OpenStudy (anonymous):

oh this one is tricky \[\frac{x}{4-x}=\frac{-x}{x-4}\]

OpenStudy (anonymous):

so now \[x+3-x=3\] and you get \[\frac{3}{x-4}\]